OpenStudy (maddie82502):

solve -3/4q=12 q=blank fill in the blank fan and medal

OpenStudy (maddie82502):

@Tyler.Evan2

OpenStudy (cathyangs):

You can perform the same function to both sides, so the two remain equal to each other. -3/4q=12 and divide both sides by some number (hint: a fraction!) to get the equation in the form q =

OpenStudy (maddie82502):

@Abstract_Undercover

Vocaloid (vocaloid):

start by multiplying both sides by 4

OpenStudy (maddie82502):

ok

Vocaloid (vocaloid):

-3/4q*4 = ?

OpenStudy (maddie82502):

-12/1 I think

Vocaloid (vocaloid):

nope

Vocaloid (vocaloid):

(-3/4)*4 = ?

Vocaloid (vocaloid):

still there?

OpenStudy (anonymous):

-16 because - 3/4 x -16 = 12

OpenStudy (anonymous):

So, -16

OpenStudy (maddie82502):

yea thanks @Tyler.Evan2 and @Vocaloid

OpenStudy (anonymous):

Np Anytime! Don't forget to close :)